OpenStudy (anonymous):

f=9/5 C + 32 solve for fahrenheit

OpenStudy (mathstudent55):

It's done.

OpenStudy (nincompoop):

have you tried?

OpenStudy (nincompoop):

lol @mathstudent55

zepdrix (zepdrix):

Bianca, the equation is already solved in terms of Fahrenheit. Do you have a specific Celcius that you need to plug in?

OpenStudy (abb0t):

Isn't it already solved for \(Fº\)?

OpenStudy (nincompoop):

http://www.celsiusconverter.com

OpenStudy (anonymous):

yes it is a manipulating literal equation

OpenStudy (goformit100):

Can Anyone please explain me how to derive the formula :- f=9/5 C + 32

OpenStudy (abb0t):

You need to ask your questions in another thread @goformit100 not post a questions within someone else's question. http://openstudy.com/code-of-conduct

OpenStudy (skullpatrol):

@goformit100 There is no "derivation" of this formula it is just used for conversion between scales.

OpenStudy (goformit100):

@Preetha is there derivation" of this formula ??

OpenStudy (anonymous):

in mean this is the problem theres a degree in f so 'f=9/5'c+32

OpenStudy (goformit100):

Ok HELP HERE... http://openstudy.com/study#/updates/520475cbe4b01dab337024fd for the derivation

OpenStudy (e.mccormick):

@ailabianca it looks like you need to learn more about Solving Literal Equations. http://www.purplemath.com/modules/solvelit.htm Basically, it is an application of the order of operations, but in reverse. You are looking to move things around until you isolate the goal. \(x=\dfrac{y}{2}+5\) solve for y. Normally, division happens first and addition last, but you are undoing things, so you cancel the addition first with a subtraction: \(x-5=\dfrac{y}{2}+5-5 \implies x-5=\dfrac{y}{2}\) Now you need to multiply through by 2. \((x-5)\cdot 2=\dfrac{y}{2}\cdot 2\implies 2x-10=y\) Your question would work the same way, but it is already solved for F.

OpenStudy (e.mccormick):

Every term is in degrees: \(F^{\circ}=9/5 C^{\circ} + 32^{\circ}\)

OpenStudy (e.mccormick):

http://www.albireo.ch/temperatureconverter/formula.htm

OpenStudy (anonymous):

thank you

OpenStudy (e.mccormick):

np. They just leave off the degrees most times because it is understood.

OpenStudy (e.mccormick):

Remember to close things when you are done. \(\large\ddot \smile\)

OpenStudy (anonymous):

x=A(B+C) solve for c

OpenStudy (e.mccormick):

OK.... so, remember what I said about you have to undo things in reverse order from normal order of operations? So what do you need to get rid of first from near C?

OpenStudy (anonymous):